OpenStudy (surryyy):

58 1/3% as a fraction

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

$$\Huge 58\frac{1}{3}\%$$

OpenStudy (skullpatrol):

$$\Huge \%=\frac{1}{100}$$

OpenStudy (skullpatrol):

Percent means "hundredths" or "divided by 100".

OpenStudy (skullpatrol):

$$\Huge x\div100=\frac{x}{100}=x\cdot\frac{1}{100}$$

OpenStudy (skullpatrol):

or x%